# LLM4SBR: A Lightweight and Effective Framework for Integrating Large Language Models in Session-based Recommendation

### v1

@@ 3. METHODOLOGY
The overall architecture of LLM4SBR is depicted in Figure 1. This section will introduce the problem definition and the specific implementation details of each module in turn.
Figure 1. LLM4SBR framework diagram. LLM4SBR is a two-stage framework. In the first stage—the Intent Inference Stage, LLM makes initial inferences based on prompts from different perspectives (long-term and short-term). Subsequently, the intent localization module is utilized to eliminate hallucinations and enhance semantics in the inference results, with the embeddings of the results stored in an external file as text data. Then, in the second stage—Representation Enhancement Stage—interaction data and text data are synchronously loaded into the model. Traditional SBR models are used to model the interaction data to obtain local and global session representations. Meanwhile, we parse the inference embeddings stored in text format for each perspective and restore them to tensor form for subsequent computations. After aligning and uniforming session representations and inference representations of the same perspective, all representations are fused into the final session representation for prediction.
3.1. Problem Formulation
The objective of SBR is to predict the next interaction item expected to occur in the current session history of an anonymous user. Here, we provide the problem definition in mathematical terms. Each data entry in the dataset represents a session sequence. Let the collection of all sessions be denoted as $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{m}\}$, where $m$ is the total number of sessions. The item set is the summary of items that have appeared in all sessions, which we define as =$\mathcal{I}=\{i_{1},i_{2},\cdots,i_{n}\}$, where $n$ is the total number of items in the set. We represent the t-th session $s_{t}$ in the dataset as $s_{t}=\{i_{t,1},i_{t,2},\cdots,i_{t,k},\cdots,i_{t,|s_{t}|}\}$, where $|s_{t}|$ is the length of the current session, and $i_{t,k}\in\mathcal{I}$ represents the $k$-th clicked item in the current session $s_{t}$. Based on the above symbols and descriptions, we define the modeling goal of session $s_{t}$ as predicting the click of the $|s_{t}|+1$th item based on the historical behavior records of $s_{t}$.
@@ 3.2.1. Prompt Design
Figure 2. Illustration of the design of prompts.
At the current stage, the logical reasoning ability of LLM is limited. To achieve more accurate inference, we introduce perspective-limiting qualifiers as auxiliary, enabling LLM to speculate on existing items in a sequence from a specific perspective rather than directly predicting using LLM.
Specifically, in our prompt design, we utilize the perspective-limiting qualifiers based on commonly used behavioral modeling perspectives in SBR (long-term and short-term). By artificially setting them, we decompose the text inference task into finer-grained perspective inference subtasks, thereby maximizing the utilization of LLM’s reasoning capabilities. It is worth noting that the perspective settings are not fixed and can be freely added or removed, endowing the framework with scalability.
The specific prompt design is illustrated in Figure 2, where we denote perspective-limiting qualifiers with blue color. A prompt consists of three parts: [background prompt, item name sequence, and task prompt ]. Some studies (Li et al., 2023c; Hua et al., 2023) suggest that ID information helps LLM distinguish between different items more accurately. Inspired by this, we incorporate corresponding ID information after the item names in the prompt design. Therefore, we present the prompt template as follows:
”The order in which users click on items is as follows:
1. ItemName__\__ItemID
$\cdots$
N. ItemName__\__ItemID
Please guess an item that the user is interested in in the (perspective-limiting qualifiers). (Only output the item name without any explanation.)”
@@ 3.2.2. LLM Inference
To enhance the effective utilization of semantic information and understand the genuine intent of sessions, we leverage the contextual understanding and logical reasoning capabilities of LLMs to achieve intent inference from different perspectives.
In the aspect of selecting large models, we have chosen the Qwen-7B  [^1] https://github.com/QwenLM/Qwen model as the inference model after comprehensive consideration of LLM’s inference capability, adaptability to both Chinese and English languages, and model parameter count. It is worth noting that here, the LLM is interchangeable. LLMs with more parameters and stronger reasoning capabilities can produce more accurate inference results. We adopt the form of question and answer, input different perspective prompts as questions to the LLM, and then the LLM returns its inferring results according to the prompts. In addition, to standardize the answers of the LLM, we specially mark ”(Only output the item name without any explanation.)” in the prompts.
@@ 3.2.3. Intent Localization
To assist LLM in eliminating hallucinations and achieving and achieving semantic enhancement, we designed the intent localization module. Although in most cases, the LLM inference result is an accurate item name, sometimes it may be just a vague item category or key project term. In rare cases, a reasonable inference result may not be obtained. The red portion in Figure 3 illustrates the initial inference results of LLM.
Inspired by the RAG retrieval model (Lewis et al., 2020), addressing hallucinations in LLM requires providing relevant external knowledge to LLM. The text retrieval scheme of the RAG model is usually based on the similarity of text embeddings, so we first encode all inference results and the text of the item set into embedding forms using a pre-trained BERT model (Devlin et al., 2018) [^2] https://huggingface.co/bert-base-uncased.
@@
where $e_{\text{infer}},e_{\text{item}}\in\mathbb{R}^{d_{text}}$.
Then, we compute the cosine similarity scores between each inference result and all item embeddings. Utilizing text embedding similarity, we select the Top-$f$ most similar actual items from the item set, where $f$ is a hyperparameter that controls the number of semantically similar items to be filtered. We multiply the embeddings of selected items by their corresponding similarity scores and then sum them up to obtain the inference result of the LLM. Figure 3 illustrates the comparison of inference results before and after using the intent localization module. Finally, the inference results from each perspective undergo hallucinations removal and semantic enhancement through this module. To reduce unnecessary computation time in the next stage, we store the adjusted results’ embeddings in an external file.
(3) $Similarity=\frac{{e_{\text{infer}}^{i}}^{\intercal}e_{\text{item}}^{j}}{\|e_{\text{infer}}^{i}\|\|e_{\text{item}}^{j}\|},$
(4) $h_{\text{infer}}=\sum_{j=1}^{f}{\frac{\exp(Similarity)}{\sum_{j=1}^{f}{\exp(Similarity)}}e_{\text{item}}^{j}},$
where $e_{\text{infer}}^{i}\in E_{\text{infer}}$ is the text embedding of the inference result and $e_{\text{item}}^{i}\in E_{\text{item}}$ is the text embedding of the item name.
@@ 3.3. Representation Enhancement Stage
After the intent inference stage, we move into the representation enhancement phase. In this stage, the SBR model processes behavioral modeling data and parsed inference data. Subsequently, the alignment and uniformity of session embeddings and inference embeddings are conducted separately for each perspective. Ultimately, all perspective inference embeddings are fused with session embeddings to form the final session representation used for prediction.
@@ 3.3.1. SBR Modeling
3.3.2. Text Embeddings Parsing
In the intent inference stage, we save the inference results’ embeddings in an external file. Therefore, we need to read out the embeddings of the inference, then parse and restore them into tensor form, followed by performing dimension alignment.
(6) $H_{\text{infer}}=W_{1}(\text{Parse}(H_{\text{infer}}^{\text{str}}))+b_{1}$
Here, $w_{1}\in\mathbb{R}^{d\times d_{text}}$ is the weight matrix, and $b_{1}\in\mathbb{R}^{d}$ is the bias term. $H_{\text{infer}}^{\text{str}}$ represents the inference embeddings stored as strings, and Parse denotes the conversion between strings and embeddings using the ”$ast.literal\_eval$” function.
@@ 3.3.3. Representation Alignment and Fusion
The SBR model models interaction information within sessions, while LLM employs its knowledge to infer the textual content corresponding to sessions. Although both have the same goal, they are not in a unified embedding space. To better integrate embeddings and enhance session representation quality, we incorporate DirectAU (Wang et al., 2022b) for alignment and uniformity of representation.
(7) $\mathcal{L}_{a}=\underset{(infer,t)\sim sess}{\mathbb{E}}||\tilde{h_{\text{infer}}^{p}}-\tilde{h_{t}^{p}}||^{2}$
(8) $\mathcal{L}_{u}={\log e^{-2{||\tilde{h_{\text{infer}}^{p}}-\tilde{h_{\text{infer}^{{}^{\prime}}}^{p}}||}^{2}}/2}+{\log e^{-2{||\tilde{h_{t}^{p}}-\tilde{h_{t^{{}^{\prime}}}^{p}}||}^{2}}/2}$
where $\mathcal{L}_{a}$ denotes alignment loss function and $\mathcal{L}_{u}$ denotes uniformity loss function. For each perspective (long-term, short-term), we separately compute the alignment loss between the inference representation and session representation under that perspective, as well as the uniform loss within each inference representation and each session representation. $h_{\text{infer}}^{p}$ and $h_{t}^{p}$ represent the inference representation and session representation, respectively, corresponding to the session $t$ under the same perspective.
Then, we fuse the session representations from different perspectives and modalities into the final session representation.
(9) $H_{sess}=W_{2}[H_{t}^{l};H_{t}^{g};H_{\text{infer}}^{\text{lt}};H_{\text{infer}}^{\text{st}}],$
where $W_{2}\in\mathbb{R}^{d\times 4d}$ is a weight matrix. $H_{t}^{l}$ is the local preference representation obtained in the SBR model, where the local preference embedding is simply defined as the last clicked item. $H_{t}^{g}$ is the global embedding obtained by the SBR model, which is obtained by the soft attention mechanism. For details, please see SR-GNN (Wu et al., 2019). Additionally, $H_{\text{infer}}^{\text{st}}$ and $H_{\text{infer}}^{\text{lt}}$ represent the short-term and long-term perspective text embeddings of LLM inference, respectively. From this, the four embeddings are compressed into the same embedding space through a linear layer.
@@ 3.3.4. Prediction and Optimization
(10) $\hat{y_{i}}=softmax(h_{sess}^{T}v_{i}),$
@@
(12) $\mathcal{L}=\mathcal{L}_{r}+\tau\left({L}_{a}+{L}_{u}\right)$
@@ 4.1.1. Datasets
We initially hoped to validate the performance using commonly used datasets in SBR, as they are more representative. Unfortunately, we couldn’t find any session datasets that provide both interaction ID sequences and item name information. Taking this into consideration, we opted for Beauty [^3] https://jmcauley.ucsd.edu/data/amazon/links.html and MovieLens-1M (ML-1M) [^4] https://grouplens.org/datasets/movielens/ datasets, which are the closest in format. The details of these two datasets are shown in Table 1. For both datasets, we adhere (Wu et al., 2019; Li et al., 2017) to removing sessions with a length of $1$ and items that appear fewer than $5$ times across all sessions.
  • Beauty dataset comprises evaluations and ratings from users on various beauty products. We treat all ratings sequences from a single user as a session sequence. We enhance the dataset using the commonly employed sequence segmentation method (Li et al., 2017; Liu et al., 2018; Wu et al., 2019) in SBR. For instance, consider an original session $s=[i_{t,1},i_{t,2},\cdots,i_{t,n}]$. After segmentation by sequence, we obtain $([i_{t,1}],i_{t,2}),([i_{t,1},i_{t,2}],i_{t,3})$, $\cdots$ ,$([i_{t,1},i_{t,2},\cdots,i_{t,n-1}],i_{t,n})$.
  • ML-1M dataset consists of over $1$ million ratings from more than $6,000$ users on over $4,000$ movies. Considering our research question, we partition the movie rating data of the same user in this dataset into multiple session sequences using a 10-minute interval as the splitting point.

### v2

@@ 4. METHODOLOGY
The overall architecture of LLM4SBR is depicted in Figure 1, and the framework process is shown in Algorithm 1. This section will introduce the problem definition and the specific implementation details of each module in turn.
Figure 1. LLM4SBR framework diagram. LLM4SBR is a two-stage framework: (a) In the intent inference stage, LLM makes initial inferences based on prompts from different views (long-term and short-term). Subsequently, the intent localization module is utilized to alleviate hallucinations and enhance semantics in the inference results. (b) In the representation enhancement stage, interaction data and text data are synchronously loaded into the model. Traditional SBR models are used to model the interaction data to obtain local and global session representations. After aligning and uniforming session representations and inference representations of the same view, all representations are fused into the final session representation for prediction.
@@ 4.1.1. Prompt Design
To more effectively utilize the inference capabilities of LLM and better integrate the inferential results of LLM with the representations learned by the SBR model, we introduce view constraint qualifiers as an auxiliary tool. Specifically, in our prompt design, we utilize the view-limiting qualifiers based on commonly used behavioral modeling views in SBR (long-term and short-term). By artificially setting them, we decompose the text inference task into finer-grained view inference subtasks, thereby maximizing the utilization of LLM’s inference capabilities. It is worth noting that the view settings are not fixed and can be freely added or removed, endowing the framework with scalability.
[…]
Figure 2. Illustration of the design of prompts. LLM will perform inference based on short-term and long-term prompts respectively to obtain inference results from different views.
@@ 4.1.2. LLM Inference
Figure 3. An example of LLM inference from different views.
To enhance the effective utilization of semantic information and understand the genuine intent of sessions, we leverage the contextual understanding and logical inference capabilities of LLMs to achieve intent inference from different views. Figure 3 shows an example of LLM reasoning based on prompts from different views. It is worth noting that the LLM is interchangeable here. LLMs with more parameters and stronger inference capabilities can produce more accurate inference results. We adopt the question-and-answer format, input different view prompts as questions to the LLM, and then the LLM returns its inferring results according to the prompts.
(1) $\text{Text}_{\text{infer}}=\textbf{\text{LLM}}(\text{prompt}_{\text{p}})$
@@ 4.1.3. Intent Localization
To assist LLM in alleviating hallucinations and achieving semantic enhancement, we designed the intent localization module. Although in most cases, the LLM inference result is an accurate item name, sometimes it may be just a vague item category or key item term. In rare cases, a reasonable inference result may not be obtained. The red portion in Figure 4 illustrates the initial inference results of LLM.
Inspired by the RAG retrieval model (Lewis et al., 2020), alleviating hallucinations in LLM requires providing relevant external knowledge to LLM. The text retrieval scheme of the RAG model is usually based on the similarity of text embeddings, so we first encode all inference results and the text of the item set into embedding forms using a pre-trained BERT model (Devlin et al., 2018) [^1] https://huggingface.co/bert-base-uncased.
@@
where $e_{\text{infer}},e_{\text{item}}\in\mathbb{R}^{d_{\text{text}}}$.
Then, we compute the cosine similarity scores between each inference result and all item embeddings. Utilizing text embedding similarity, we select the Top-$r$ most similar actual items from the item set, where $r$ is a hyperparameter that controls the number of semantically similar items to be filtered. We multiply the embeddings of selected items by their corresponding similarity scores and then sum them up to obtain the inference result of the LLM. Figure 4 illustrates the comparison of inference results before and after using the intent localization module. Finally, the inference results of each view are passed through this module to alleviate hallucinations and enhance semantics.
(4) $\text{Similarity}^{i}=\frac{{e_{\text{infer}}^{i}}\cdot e_{\text{item}}^{j}}{\|e_{\text{infer}}^{i}\|\|e_{\text{item}}^{j}\|},$
(5) $h_{\text{infer}}=\sum_{i_{r}\in I_{r}}\text{Similarity}^{i_{r}}e_{\text{item}}^{i_{r}},$
where $e_{\text{infer}}^{i}\in E_{\text{infer}}$ is the text embedding of the inference result and $e_{\text{item}}^{j}\in E_{\text{item}}$ is the text embedding of the item name. The formula (4) represents the calculation of the cosine similarity scores between each inference result and all item names. $I_{r}$ denotes the set of the Top-$r$ item indices with the highest similarity scores, in which $r$ is a hyperparameter.
Algorithm 1 Process details of the LLM4SBR framework.
0:  prompts(Short &\&& Long views), $I_{t}$, $\text{Text}_{\text{item}},$
0:  $\hat{y_{i}}$
1:  $E_{\text{item}}=\textbf{BERT}(\text{Text}_{\text{item}})$
2:  for each prompt $p\in\{P_{short},P_{long}\}$ do
3:     $\text{Text}_{\text{infer}}=\textbf{LLM}(p)$
4:     $E_{\text{infer}}=\textbf{BERT}(\text{Text}_{\text{infer}})$
5:     $\text{Similarity}^{i}=\frac{{e_{\text{infer}}^{i}}\cdot e_{\text{item}}^{j}}{\|e_{\text{infer}}^{i}\|\|e_{\text{item}}^{j}\|}$
6:     $h_{\text{infer}}=\sum_{i_{r}\in I_{r}}\text{Similarity}^{i_{r}}e_{\text{item}}^{i_{r}}$, where $I_{r}$ is the set of the most similar $r$ items.
7:  end for
8:  $H_{t}^{l},H_{t}^{g}=\textbf{\text{SBR}{-}\text{Model}}(\mathcal{I}_{t})$
9:  Initial $\mathcal{L}_{a}=0,\mathcal{L}_{u}=0$
10:  for each views p (long-term & short-term) do
11:     $\tilde{h_{\text{infer}}^{\text{p}}}=\mathbf{T}h_{\text{infer}}^{\text{p}}$
12:     $\mathcal{L}_{a}^{\text{p}}=\underset{(\text{infer},\text{t})\sim sess}{\mathbb{E}}||\tilde{h_{\text{infer}}^{\text{p}}}-h_{\text{t}}^{\text{p}}||^{2}$
13:     $\mathcal{L}_{u}^{\text{p}}={\log e^{-2{||\tilde{h_{\text{infer}}^{\text{p}}}-\tilde{h_{\text{infer}^{{}^{\prime}}}^{\text{p}}}||}^{2}}/2}+{\log e^{-2{||h_{t}^{\text{p}}-h_{t^{{}^{\prime}}}^{\text{p}}||}^{2}}/2}$
14:     $\mathcal{L}_{a}+=\mathcal{L}_{a}^{\text{p}}$, $\mathcal{L}_{u}+=\mathcal{L}_{u}^{\text{p}}$
15:  end for
16:  $\mathcal{L}_{a}=\bar{\mathcal{L}_{a}},\mathcal{L}_{u}=\bar{\mathcal{L}_{u}}$
17:  $\alpha_{\text{infer}}^{\text{lt}}=\textbf{Q}_{1}^{\mathrm{T}}(\text{sigmoid}(H_{\text{t}}^{g}+H_{\text{infer}}^{\text{lt}}))$
18:  $\alpha_{\text{infer}}^{\text{st}}=\textbf{Q}_{2}^{\mathrm{T}}(\text{sigmoid}(H_{\text{t}}^{l}+H_{\text{infer}}^{\text{st}}))$
19:  $H_{\text{sess}}=\textbf{W}[H_{\text{t}}^{l}*\alpha_{\text{infer}}^{\text{st}};H_{\text{t}}^{g}*\alpha_{\text{infer}}^{\text{lt}}],$
20:  $\hat{y_{i}}=\text{softmax}(h_{\text{sess}}^{\mathrm{T}}v_{i})$
21:  $\mathcal{L}_{r}=-\sum_{i=1}^{n}y_{i}log(\hat{y_{i}})+(1-y_{i})log(1-\hat{y_{i}})$
22:  $\mathcal{L}=\mathcal{L}_{r}+\tau\left({L}_{a}+{L}_{u}\right)$
23:  return  $\hat{y_{i}}$
@@ 4.2. Representation Enhancement Stage
After the intent inference stage, we move into the representation enhancement phase. In this stage, the SBR model processes behavioral modeling data and parsed inference data. Subsequently, the alignment and uniformity of session embeddings and inference embeddings are conducted separately for each view. Ultimately, all view inference embeddings are fused with session embeddings to form the final session representation used for prediction.
@@ 4.2.2. Representation Alignment and Fusion
The SBR model models the interaction information in a session, while the LLM uses its knowledge to infer the semantic content corresponding to the session. Although both have the same goal, they are not in a unified embedding space. Here we draw on the approach of DirectAU (Wang et al., 2022b) to unify the SBR and LLM representations into a common space to achieve alignment and unification of the representations in different viewpoints.
(7) $\tilde{h_{\text{infer}}^{\text{p}}}=\mathbf{T}h_{\text{infer}}^{\text{p}},$
(8) $\mathcal{L}_{a}=\underset{(\text{infer},\text{t})\sim sess}{\mathbb{E}}||\tilde{h_{\text{infer}}^{\text{p}}}-h_{\text{t}}^{\text{p}}||^{2},$
(9) $\mathcal{L}_{u}={\log e^{-2{||\tilde{h_{\text{infer}}^{\text{p}}}-\tilde{h_{\text{infer}^{{}^{\prime}}}^{\text{p}}}||}^{2}}/2}+{\log e^{-2{||h_{t}^{\text{p}}-h_{t^{{}^{\prime}}}^{\text{p}}||}^{2}}/2},$
where the matrix $\mathbf{T}\in\mathbb{R}^{d\times d_{\text{text}}}$ transforms the embedding vectors of the text modal into the latent space $\mathbb{R}^{d}$, $\mathcal{L}_{a}$ denotes alignment loss function and $\mathcal{L}_{u}$ denotes uniformity loss function. For each view (long-term, short-term), we separately compute the alignment loss between the inference representation and session representation under that view and the uniform loss within each inference representation and each session representation.
We then fuse session representations from different views and modalities into a final session representation through a soft-attention mechanism.
(10) $\alpha_{\text{infer}}^{\text{lt}}=\textbf{Q}_{1}^{\mathrm{T}}(\text{sigmoid}(H_{\text{t}}^{g}+H_{\text{infer}}^{\text{lt}})),$
(11) $\alpha_{\text{infer}}^{\text{st}}=\textbf{Q}_{2}^{\mathrm{T}}(\text{sigmoid}(H_{\text{t}}^{l}+H_{\text{infer}}^{\text{st}})),$
(12) $H_{\text{sess}}=\textbf{W}[H_{\text{t}}^{l}*\alpha_{\text{infer}}^{\text{st}};H_{\text{t}}^{g}*\alpha_{\text{infer}}^{\text{lt}}],$
where parameters $\textbf{Q}_{1},\textbf{Q}_{2}\in\mathbb{R}^{d}$ and $\textbf{W}\in\mathbb{R}^{d\times 2d}$ transforms the concatenated embedding vectors into a latent space $\mathbb{R}^{d}$. $H_{\text{t}}^{l}$ is the local preference representation obtained in the SBR model, where the local preference embedding is simply defined as the last clicked item. $H_{\text{t}}^{g}$ is the global embedding obtained by the SBR model, which is obtained by the soft attention mechanism. For details, please see SR-GNN (Wu et al., 2019). Additionally, $H_{\text{infer}}^{\text{st}}$ and $H_{\text{infer}}^{\text{lt}}$ represent the short-term and long-term view text embeddings of LLM inference, respectively. Finally, we concatenate the weighted local and global representations into the final session representation $H_{\text{sess}}$.
@@ 4.2.3. Prediction and Optimization
(13) $\hat{y_{i}}=\text{softmax}(h_{\text{sess}}^{\mathrm{T}}v_{i}),$
@@
(15) $\mathcal{L}=\mathcal{L}_{r}+\tau\left(\mathcal{L}_{a}+\mathcal{L}_{u}\right),$
@@
4.3. Discussion
Ease-of-Use: One key aspect that makes LLM4SBR easy to use is its modular design. The separation of Intent Inference and Representation Enhancement stages allows developers to integrate or modify individual components without necessarily disrupting the entire system. Moreover, LLM4SBR retains the data interface of the traditional SBR model, which means that you only need to prepare the data required for LLM inference to seamlessly transition to the training phase without making large-scale changes to the existing workflow. This makes the migration from the old system to the new model smoother and more efficient.
Plug-and-Play Compatibility: The LLM4SBR framework utilizes SBR models to process interaction data, ensuring that advances in session modeling can be easily adopted without the need for radical modifications. This compatibility encourages researchers to experiment with different SBR models as “plug-ins,” choosing the most appropriate model based on their specific application requirements or performance benchmarks. Additionally, the LLM in the LLM4SBR framework can also be updated and does not require dedicated pre-training and fine-tuning. As research in the LLMs continues to advance, the LLMs in the framework can be replaced by newer and more powerful models. This feature enables it to adapt to the ever-changing research environment and technological advances.
@@ 5. Experiments
In this section, we design a series of experiments to answer the following four questions:
  • RQ1: Can the LLM4SBR framework improve the performance of the SBR model?
  • RQ2: Is each component in the LLM4SBR framework necessary and what is its impact on the overall performance?
  • RQ3: How do hyperparameter settings affect performance in LLM4SBR?
  • RQ4: Is the LLMSBR framework lightweight?
@@ 5.1.1. Datasets
We initially hoped to use datasets commonly used in SBR, such as Diginetica, Nowplaying, and RetailRocket, etc., to validate performance as they are more representative. Unfortunately, none of these datasets provide both interaction ID sequences and item name information. Consequently, we selected the classic datasets Beauty and MovieLens-1M (Ml-1M) in the Sequential Recommendation, and adapted them to a session format.
The details of these two datasets are shown in Table 2. For both datasets, we adhere (Wu et al., 2019; Li et al., 2017) to removing sessions with a length of $1$ and items that appear fewer than $5$ times across all sessions.
  • Beauty  [^2] https://jmcauley.ucsd.edu/data/amazon/links.html dataset comprises evaluations and ratings from users on various beauty items. We treat all ratings sequences from a single user as a session sequence. We enhance the dataset using the commonly employed sequence segmentation method (Li et al., 2017; Liu et al., 2018; Wu et al., 2019) in SBR. For instance, consider an original session $s=[i_{t,1},i_{t,2},\cdots,i_{t,n}]$. After segmentation by sequence, we obtain $([i_{t,1}],i_{t,2}),([i_{t,1},i_{t,2}],i_{t,3})$, $\cdots$ ,$([i_{t,1},i_{t,2},\cdots,i_{t,n-1}],i_{t,n})$.
  • Ml-1M  [^3] https://grouplens.org/datasets/movielens/ dataset consists of over $1$ million ratings from more than $6,000$ users on over $4,000$ movies. Considering that our research problem is SBR, we have performed special processing on the sequence dataset by using 10-minute intervals as segmentation points to divide the user sequence into multiple session sequences.